multiply by 9 if you add the answers digits together, you get 9. Example: 9×5=45 and 4+5=9. (But not with 9×11=99)

multiply by 9


if you add the answers digits together, you get 9. 
Example: 9×5=45 and 4+5=9. (But not with 9×11=99)

Grade:

2 Answers

Nishant Vora IIT Patna
9 years ago
Hello Student,

Please post your question properly.
Thanks,
______________________________________________________________________________________________________________________________________________________________________________________
Yugabrat Gogoi
9 years ago
ROHAN,
 
ALL THE MULTIPLE OF 9 from 1 to infinite . LIKE-9 X 1=09 and 0+9=9
9 x 2=18 and 1+8=9 in this way 9x11=99 and 9+9=18 and here 1+8=again 9.
9x1=9 and 0+9=9
9x2=18 and 1+8=9
9x3=27 and 2+7=9........
and so on like 999999999999999999999999999999999999999999 X 1= 999999999999999999999999999999999999999999 and when we plus 9+9+9+9and so on and we plus the result till its a single digit and we see a magic, its 9.